# Towards Automated Semantic Interpretability in Reinforcement Learning via Vision-Language Models

### v2

@@ 3.2.1 Interpretable control tree
Interpretable Control Tree (ICT) is adapted from Interpretable Continuous Control Tree (ICCT) proposed by Paleja et al.  [18]. The ICT is structured similarly as ICCT, where each decision node determines the routing of input features, $\mathbf{x}$, to sparse linear controllers at the leaf nodes.
@@
Since the $\arg\max$ operator is non-differentiable, we approximate this function with the Gumbel-Softmax  [11], as per Eq. 5, where $g_{j}\sim\text{Gumbel}(0,1)$ and $\tau$ is a temperature parameter.
@@
During inference, a hard threshold is applied to enforce discrete decision-making using only one feature per node, as per Eq. 6.
@@
While during training, gradients are allowed to flow through discrete decisions using the straight-through trick [3]. This approach ensures that each decision node remains interpretable while allowing for end-to-end gradient-based learning.
@@
Unlike ICCT, ICT supports both discrete and continuous actions without re-architecture, which helps generalizability. In the discrete case, leaf nodes in ICT output the possibility of taking each action as a Probability Mass Function (PMF). If there are six possible actions in the discrete action space, then there will be six sparse linear functions of the input features followed by a softmax function at each leaf node. And for the continuous action space, leaf nodes output the Probability Density Function (PDF) over possible actions.
4 Experiments
4.1 Environment
@@
We evaluate our approach across three diverse environments: Atari Skiing, a fast-paced pixel-art game requiring gate navigation; BabyAI-GoToRedBall, a symbolic grid world for instruction following; and OpenAI Gym Highway, a 2D environment for high-level autonomous driving decisions like lane switching and speed control. The Atari Skiing environment provides raw image state spaces without any pre-defined feature vectors or feature sets. BabyAI and Highway offer default feature vectors, but these vectors describe the full state of the environment rather than focusing on the task-relevant semantic features needed for decision making. By selecting these environments, we ensure that the VLM engages in genuine reasoning over the environment to identify key semantic features, rather than relying on prior internet-based knowledge, a phenomenon referred to as training data memorization [5].
4.2 Feature extraction
4.2.1 VLM deducing key semantic features.
@@ 4.2 Feature extraction
For this step, we utilize ChatGPT-4o-latest, one of the current leading models in language reasoning. Across all three environments, the VLM consistently produces a deterministic set of semantic features for decision-making, achieving performance comparable to human reasoning. For all three environments, we query the VLM 50 times and use majority vote to get the most prominent answer from the VLM. A complete set of those features are included in Appendix A.2.2.
4.2.2 VLM generating Feature Extraction code.
@@
For this step, we utilize claude-3.5-sonnet-20241022, the current leading model in code generation. To address common challenges with VLM-generated code, such as non-functionality or runtime errors, we query the VLM 50 times using the same prompt (temperature = 0 for reproducibility) and select the most reliable feature extraction code through majority voting.
@@
The generated codes are evaluated on a randomly collected dataset of 100 images and filtered based on the following criteria: (1) the code must compile and execute successfully and (2) it must produce diverse feature outputs rather than constant or repetitive values across the dataset. Candidates producing identical outputs are grouped into clusters, and the largest cluster is chosen as the final feature extraction code. Details are included in Appendix A.2.
4.2.3 Extractor module performance analysis
In Step 3 of our methodology, we address the inefficiencies of the VLM-generated feature extraction code for single-image frames, $\mathcal{G}_{s}$, by training an lightweight Extractor Module as its replacement. While $\mathcal{G}_{s}$ can be executed directly on the dataset, the processing pipeline is inherently non-batchable, as it handles image segmentations independently. This non-batchability severely impacts RL training, causing slowdowns of around 20000× compared to batchable operations (our CNN Extractor Module), as demonstrated in Table 3. From practical experience, this would translate to training times exceeding multiple weeks on an H40-level GPU for a single policy to converge.
@@ 4.2.3 Extractor module performance analysis
We constrain computing time to a reasonable range with a lightweight Extractor Module, resolving the slowdown caused by segmentation-based analysis in $\mathcal{G}_{s}$. For each environment, we train an CNN on 7,000 images collected from diverse trajectories, validate it on 2,000 images, and test it on 1,000 images. Compared to the VLM-generated ground truth, our models achieve high accuracy (typically >0.95) on most binary features, while MSE remains low (<0.005) for the majority of continuous features, with one exception where error is notably higher. Table 1 provides detailed performance statistics. These results confirm that a lightweight vision model can serve as a reliable and computationally efficient replication for VLM generated extraction code $\mathcal{G}_{s}$.
@@
Table 1: Performance of the Extractor Module across environments, compared with ground truth from VLM-generated code. Continuous components are evaluated with MSE; boolean components are evaluated with accuracy. Dashes indicate vacant entries due to varying feature lengths and types across environments. Full feature descriptions are provided in Appendix A.2.2.
@@ 4.3 Training results
  • CNN: An uninterpretable black-box model that maps raw visual inputs directly to the action space, trained with RL.
@@
Table 2: Runtime speedup on Skiing.
Other domains in Appendix A.3.
Batch VLM Code CNN Speedup
8 19.056 0.001 13831.3x
16 38.288 0.002 17325.6x
32 76.212 0.004 19935.0x
64 151.836 0.007 21692.9x
128 308.468 0.014 22772.4x
Figure 3: Results across three domains. Higher scores are better. Runtime comparison shows speedup of CNN-based extractor over VLM code.
@@
  • VLM-DT (claude-3.5-sonnet): An equivalent baseline using Claude 3.5 Sonnet as the prompting engine.
@@
We use Proximal Policy Optimization (PPO) as the reinforcement learning algorithm to train our policy [24], with a learning rate of 0.0005 and a discount factor of 0.99. During testing, the policy is tested five times using different random seeds, each distinct from those used during training and evaluation. Results are shown in Fig. 3. Across all three environments, VLM-ICT consistently outperforms other uninterpretable models (MLP) built on the semantic feature space, and demonstrates performance close to the CNN+RL black-box benchmark representing the upper limit of task performance. These consistent results highlight the robustness and effectiveness of our proposed method.
@@
Building on a semantically meaningful feature set, our ICTs contains nodes directly corresponding to human-understandable concepts. While we do not evaluate interpretability directly, this approach is consistent with prior work where performance serves as a proxy for meaningful representations [14, 36]. A previous human study [18] also supports the interpretability of such trees. We visualize the learned ICT in Appendix A.1, along with example cases demonstrating how the ICT yields plausible local explanations in example scenarios.
5 Limitations and future works
@@
Our framework is yet limited by the assumption that SAM performs reasonably well across environments, which may hinder generalizability in domains with poor segmentation. Future work could explore RL-based backpropagation to fine-tune VLMs, allowing them to iteratively improve outputs based on agent feedback [1]. This approach has the potential to enable VLMs to identify discrepancies between the underlying physics simulation and its visual representation, leading to improved feature extraction for downstream RL tasks.
6 Conclusion

### v3

@@ IV-A Interpretable control tree
Each decision node in ICT determines the routing of input features, $\mathbf{x}$, to sparse linear controllers at the leaf nodes.
@@
Since the $\arg\max$ operator is non-differentiable, we approximate this function with the Gumbel-Softmax  [36], as per Eq. 4, where $g_{j}\sim\text{Gumbel}(0,1)$ and $\tau$ is the temperature.
@@
During inference, a hard threshold is applied to enforce discrete decision-making using only one feature per node, as per Eq. 5, where $\mathbf{1}$ is the indicator function.
@@
During training, gradients flow through discrete decisions using the straight-through trick [37]. This approach ensures each decision node remains interpretable while allowing end-to-end gradient-based learning.
@@
V Experiments
V-A Environment
@@
We evaluate iTRACE in three diverse environments chosen to ensure the VLM performs genuine reasoning rather than relying on memorized knowledge [38]: Atari Skiing, a pixel-art game; BabyAI-GoToRedBall, a symbolic grid world for navigation; and OpenAI Gym Highway, a 2D driving environment. While BabyAI and Highway provide default feature vectors, these describe the full environment state rather than the task-relevant semantic features required for decision making. The Atari Skiing environment provides raw image state spaces without any pre-defined feature vectors or feature sets. BabyAI and Highway offer default feature vectors, but these vectors describe the full state of the environment rather than focusing on the task-relevant semantic features needed for decision making.
V-B Feature extraction
V-B1 VLM deducing key semantic features.
@@ V-B Feature extraction
For this step, we utilize ChatGPT-4o-latest, one of the current leading models in language reasoning. Across all three environments, the VLM consistently produces a deterministic set of semantic features for decision-making, achieving performance comparable to human reasoning. For all three environments, we query the VLM 50 times and use majority vote to get the most prominent answer from the VLM. A complete set of those features are included in Table II.
TABLE II: Feature schemas for BabyAI RedBall, Highway, and Skiing environments. Type: C = Continuous, B = Binary. Temporal: T = True, F = False.
Feature Type Temp.
BabyAI
red_ball.{x,y}_position C F
red_ball.is_visible B F
[…]
Highway
agent.{x,y}_position C F
agent.x_velocity C T
lane_existence.{upper,lower} B F
vehicle_ahead.{x_position,x_velocity} C F/T
vehicle_upper_lane.{x_position,x_velocity} C F/T
vehicle_lower_lane.{x_position,x_velocity} C F/T
Skiing
[…]
next_pole.y_position C F
next_tree.{x,y}_position C F
skier.x_position C F
skier.x_velocity C T
V-B2 VLM generating Feature Extraction code.
@@
For this step, we utilize claude-3.5-sonnet-20241022, one of the leading models in code generation. To address common challenges with VLM-generated code, such as non-functionality or runtime errors, we query the VLM 50 times using the same prompt (temperature = 0 for reproducibility) and select the most reliable feature extraction code through majority voting.
Since code generation depends on segmentation quality $\mathcal{S}$, we further evaluate segmentation performance. Among the evaluated environments, Skiing presented the greatest difficulty for VLM-based segmentation labeling, and we therefore selected it for detailed analysis. On 100 sampled images, segmentation achieves high overall accuracy (Table III) but includes imperfect cases where the skier is partially or entirely missed (Fig. 4). Despite such errors present in our in-context-learning examples (Fig. 4), the generated code remains effective, demonstrating the robustness of our framework as long as a subset of segments is correctly labeled.
(a) Skiing Exp. 1.
(b) Skiing Exp. 2.
(c) Skiing Exp. 3.
Figure 4: Randomly selected segmentation examples we provide to the VLM for domain Skiing. Example 1 is good while 2 and 3 are imperfect. Example 2 identifies the skier as a partial skier and a noisy segment as skier. Example 3 identifies both the actual skier on the right and decoration elements at the bottom as partial skier.
@@
The generated codes are evaluated on a randomly collected dataset of 100 images and filtered based on the following criteria: (1) the code must compile and execute successfully and (2) it must produce diverse feature outputs rather than constant or repetitive values across the dataset. Candidates producing identical outputs are grouped into clusters, and the largest cluster is chosen as the final feature extraction code.
TABLE III: Confusion ratios per feature concept
Skier Pole Tree
[…]
V-B3 Extractor module performance analysis
@@
We train an CNN based Extractor Module to resolve the slowdown caused by segmentation-based analysis. For each environment, we train an CNN on 7,000 images collected from diverse trajectories, validate it on 2,000 images, and test it on 1,000 images. As shown in Table IV, compared to the VLM-generated ground truth, our models achieve high accuracy (typically >0.95) on most binary features, while MSE remains low (<0.005) for the majority of continuous features, with one exception where error is notably higher. These results confirm that a lightweight vision model can serve as a reliable and computationally efficient replication for VLM generated extraction code $\mathcal{G}_{s}$.
@@
TABLE IV: Performance of the Extractor Module across environments, compared with ground truth from VLM-generated code. Continuous components are evaluated with MSE; boolean components are evaluated with accuracy. Dashes indicate vacant entries due to varying feature lengths and types across environments. Full feature descriptions are provided in Table II.
@@ V-C Baselines
  • CNN: An uninterpretable black-box model that maps raw visual inputs to the action space, trained with RL.
@@
  • VLM-EQL: An interpretable neural network that employs simple mathematical functions (e.g., square, cube, sine) to map semantic features to the action space [39].
@@
Figure 5: Results across three domains. Higher scores are better.
@@
  • VLM-DT (claude-3.5-sonnet): An equivalent baseline querying Claude 3.5 Sonnet.
V-D Train results
@@
We use Proximal Policy Optimization (PPO) as the reinforcement learning algorithm to train our policy [40], with a learning rate of 0.0005 and a discount factor of 0.99. During testing, the policy is tested five times using different random seeds, each distinct from those used during training and evaluation.
Results are shown in Fig. 5. Our results demonstrate the remarkable effectiveness of the proposed framework. As expected, the CNN policy achieves the highest performance across all three environments, benefiting from its status as a black-box model with orders of magnitude more parameters than the interpretable baselines. Notably, our proposed VLM-ICT method consistently reduces this gap. In Atari Skiing, VLM-ICT matches both the interpretable baseline (VLM-EQL) and the black-box policy (VLM-MLP) trained on the same features. In BabyAI and Highway, VLM-ICT outperforms the interpretable baseline (VLM-EQL) and reaches performance comparable to the black-box model (CNN, VLM-MLP). Furthermore, across all domains, our interpretable framework steadily outperforms decision trees generated directly by VLMs (VLM-DT ChatGPT and VLM-DT Claude), demonstrating that these are non-trivial control problems even when the semantic features provided are precisely task-relevant, with no redundant or irrelevant information.
@@
Building on a semantically meaningful feature set, our ICTs contains nodes directly corresponding to human-understandable concepts. While we do not evaluate interpretability directly, this approach is consistent with prior work where performance serves as a proxy for meaningful representations [41]. A previous human study [5] also supports the interpretability of such trees.
V-E Takeaways
Our results highlight followin achievements of the proposed iTRACE framework:
  • Fast: By distilling VLM-generated feature extractors into a lightweight CNN, iTRACE avoids repeatedly querying the segmentation model at every timestep. This design yields several orders of magnitude speedup, making RL training practical at scale.
  • Trainable: Both the feature extractor and the tree-based policy adapt effectively to new domains via gradient-based optimization. This enables pixels-to-action control and shows that interpretability does not come at the cost of flexibility.
  • Interpretable: iTRACE provides transparency at every stage. The feature space is grounded in human-understandable concepts, and the learned Interpretable Control Tree ensures the resulting policy is verifiable.
In summary iTRACE delivers a pixels-to-actions system that is simultaneously fast, trainable, and interpretable.
VI Limitations and future works
@@
Our framework is limited by two factors: first, it assumes that SAM performs reasonably well across environments, which may hinder generalizability in domains with poor segmentation; second, it requires domain knowledge, which still involves human input to some extent and is left to be automated by foundation model queries or agentic exploration. Future work could explore RL-based backpropagation to fine-tune VLMs, allowing them to iteratively improve outputs based on agent feedback [42]. This approach has the potential to enable VLMs to identify discrepancies between the underlying physics simulation and its visual representation, leading to improved feature extraction for downstream RL tasks.
VII Conclusion